OpenStudy (anonymous):

Add 5x to both sides. y2 − 5x + 5x =_____+1

OpenStudy (anonymous):

y2 − 5x + 5x =_____+1 y2 +5x = 5x + 1 is that it ?

OpenStudy (anonymous):

oops i left off the original problem y2 – 5x = 1

OpenStudy (anonymous):

are you trying to solve for Y ?

OpenStudy (anonymous):

y2 - 5x = 1 (add 5x both sides) y2 -5x +5x = 1 + 5x y2 = 1 +5x

OpenStudy (anonymous):

I have to simplify y^2 and then Use the square root property. y = ± 5